OpenStudy (anonymous):

(z-5/z+5)=3/7 Solve for z

OpenStudy (anonymous):

z = 25/2 = 12.5

OpenStudy (anonymous):

z-5=3/7(z+5) multiply both sides by (z+5) z-5= 3/7z+15/7 distribute 3/7 through z-3/7z=15/7+5 add 5 to both sides and subtract 3/7z from both sides 7/7z-3/7z=15/7 + 35/7 find a common denominator prior to adding 4/7z = 50/7 perform addition and subtraction (7) 4/7z = 50/7 (7) multiply both sides by 7 to get rid of the 7 in the denominator 4z = 50 4z/4 = 50/4 divide both sides by 4 to get z alone z=50/4 or 25/2 or 12.5